If you pile oranges in a pyramid that has 1 orange in the top layer, 4 in the next layer nine in the third layer, and 16 in the fourth layer, how many oranges will you need to make a pile with 10th layer?
Found 3 solutions by ewatrrr, stanbon, richwmiller:
Answer by ewatrrr(24785)
Answer by stanbon(75887)

If you pile oranges in a pyramid that has 1 orange in the top layer, 4 in the next layer nine in the third layer, and 16 in the fourth layer, how many oranges will you need to make a pile with 10th layer?
--------------------------------
Sum = 1 + 4 + 9 + 16 + ... + 100
-----
Sum = (n(n+1)(2n+1)/6
------
Ans = [10*11*21)/6 = 770
--------------------------------
Cheers,
Stan H.
--------------------

Answer by richwmiller(17219)

This is neither arithmetic nor geometric progression
Tn=n^2
1,4,9,16,25,36,49,64,81,100
The tenth layer will have 100 oranges
for a total of 385 oranges in the 10 layer display